# LLM-based Few-Shot Early Rumor Detection with Imitation Agent

### v1

@@ 5.1. Experiments Setup
Four real-world datasets are used: 1) PHEME (zubiaga2016learning): This dataset collects posts from conversation threads during newsworthy events; 2) TWITTER (ma2016detecting): This dataset provides posts of events sourced from a fact-checking website, widely used for rumor detection studies; 3) Twitter-COVID-19 (lin-etal-2022-detect): This dataset gathers posts from conversation threads discussing events about the COVID-19 pandemic (DBLP:journals/corr/abs-2010-06906); 4) BEARD (zeng-gao-2022-early): this dataset is for EARD task, focusing on covering early-stage posts relevant to various events. Each dataset offers a chronological collection of relevant posts for each event, annotated as either rumor or non-rumor.
@@
We compare with existing full-shot EARD methods under few-shot setup, including ERD (zhou2019early) that uses a Deep Q-Network (DQN) to enforce the model to focus on early posts; CED (song2019ced) that uses a fixed probability threshold to check if the prediction is credible for determining early detection point; HEARD (zeng-gao-2022-early) that automatically determines a detection point at which predictions in future time steps remain unchanged using the neural Hawkes Process (mei-etal-2017-The). Given the strong capability of LLMs across NLP tasks, we also include Mistral (Mistral), Llama3 (llama3) [^3] We refer to Mistral-Instruct-v0.3 as Mistral and Llama-3-Instruct as Llama3 for brevity., ChatGPT (ChatGPT) as baselines. We incorporate these LLMs using two early detection strategies. Earliest Rumor Detection Strategy: The LLMs are prompted to make predictions using only the first post of each instance, following miao2021syntax. Preset Time Checkpoints Strategy: we use a set of preset time checkpoints at $\{1\text{h},6\text{h},12\text{h},24\text{h},36\text{h}\}$. All posts published before each checkpoint are provided to the LLMs for prediction. The timestamp of the first post serves as the starting point, and the time elapsed for each subsequent post is calculated relative to this initial timestamp. This strategy is commonly used in general rumor detection studies for evaluating early detection performance (ma2016detecting; yu2017convolutional; ma2017detect; ma-etal-2018-rumor; guo2018rumor; bian2020rumor; Lin_Yi_Ma_Jiang_Luo_Shi_Liu_2023).
@@
We randomly sample 50 labeled training instances for trajectory generation and 100 unlabeled instances to serve as the environment for the agent, allowing the agent to dynamically explore and adapt its strategy in an online policy learning setting. We utilize the pretrained BERTweet model (nguyen-etal-2020-bertweet) to obtain embeddings of each state. Following ho2016generative, we employ two hidden layers with a Tanh activation function for the policy, value, and discriminator networks. The coefficient of causal entropy $\lambda$ is set to 0.01. We set the expert ratios to $\alpha=0.7$, $\beta=0.15$, and $\gamma=0.15$ as discussed in §4.2.2.
@@
We report macro-F1 score for classification and use Early Rate (ER) (song2019ced) to measure the proportion of posts used for detection. ER measures the proportion of posts used for detection: $ER=\frac{1}{|\mathbf{C}_{\text{test}}|}\sum_{C\in\mathbf{C}}\frac{i_{C}}{|C|}$, where $\mathbf{C}_{\text{test}}$ represents the test set, $i_{C}$ indicates the early detection decision made at the $i$-th post in instance $C$, and $|C|$ is the total number of posts in that instance. Lower ER means that the model can detect rumors earlier. For a fair comparison, we use the same prompt template for all models and run each experiment five times with different random seeds, reporting the mean and standard deviation for each metric. The prompt used for LLM prediction is “Analyze the given sequence of social media posts, determine if it is a rumor. Respond Yes or No only. Posts: [the list of observed posts in chronological order]”.
Table 3 presents hyper-parameters in our experiments. Following ho2016generative, the policy, value, and discriminator networks are initialized randomly. We set the number of training steps to 200,000, with a batch size of 64 for the discriminator network and 4 for the policy and value networks.
@@ 5.2. Results and Analysis
First, as shown in Table 1, despite using only first post following the earliest rumor detection strategy (miao2021syntax), LLMs achieve competitive or superior macro-F1 scores across multiple datasets, compared to existing EARD methods. Specifically, ChatGPT outperforms all existing EARD methods by up to 33.3% and 10.3% on PHEME and Twitter-COVID-19, respectively. Even smaller LLMs like Mistral surpass all existing EARD methods on PHEME, and Llama3 outperforms ERD on TWITTER dataset. These results highlight the potential of LLMs used as detectors in EARD, particularly in data-scarce scenarios.
@@
Third, our method achieves overall better performance across both ER and F1, compared to existing EARD methods. For instance, in Table 1, on the TWITTER dataset, our framework enables Mistral to surpass all EARD baselines in F1 score while achieving a significantly lower ER (0.043) compared to ERD (1.000), CED (0.617), and HEARD (0.647). We also observe that ERD fails to make early decisions in few-shot setting, as shown by its high ER scores. This is likely due to its reliance on a manually designed reward function, where its policy trained by DQN is highly sensitive. Specifically, the reward function applies only a small penalty for continuation but imposes a large penalty for incorrect early termination (zhou-etal-2019-early), discouraging the model to stop early. This behavior aligns with previous findings (zeng-gao-2022-early), and the problem is further exacerbated in the few-shot scenario, where data scarcity undermines the stability and effectiveness of reinforcement learning.
@@ 5.2.3. Analysis on How Generalizability Varies Across LLMs
As shown in Table 2 and discussed in § 5.2.1, our method generalizes better than the baseline EARD approaches to new event data. Trained on pre-COVID datasets (PHEME and TWITTER) and tested on Twitter-COVID-19, it consistently outperforms baselines in both F1 and ER. We further investigate how this generalizability varies across different base LLMs. As shown in Figure 4, our method dramatically boosts Llama3 by around 17% and 18% when trained on PHEME and TWITTER, respectively. This suggests that the agent can learn common expert decision patterns encoded in expert trajectories from other datasets, which benefits LLMs when applied to new event data. These gains are less pronounced with Mistral, which outperforms Llama3, and turn into a negative influence with ChatGPT, which initially shows the best performance on this COVID-related dataset. We conjecture that as the base model’s initial performance improves, the benefits of learned common behavior patterns become less useful compared to the information specific to the new event data. Thus, applying an agent trained on other datasets could lead to suboptimal decisions. One possible solution is to train the agent on a small number of new event data. As shown in Figure 4, even a few samples from the Twitter-COVID-19 dataset can improve performance across all base LLMs. Another promising direction is to model this problem as a Partially Observable Markov Decision Process (POMDP) (sondik1971optimal; spaan2012partially), which is useful for capturing latent uncertainties and incomplete observations in sequential decision-making (kaelbling1998planning) and may better capture the uncertainties in scenarios involving new event data. We leave this for future work.
@@ Appendix A Theoretical Analysis
The proof involves showing that the optimal policy in an imitation learning setting, typically obtained by first solving the Inverse Reinforcement Learning (IRL) problem to derive the optimal reward function $r^{*}$ and then applying a Reinforcement Learning (RL) algorithm, can be compressed into optimizing a $\psi$-regularized objective. Our method is based on generative adversarial imitation learning (ho2016generative), extending it to handle different types of expert trajectories in the context of LLM.
In the aforementioned equation, the objective is to find the saddle point of the minimax problem. Since the reward function $r(s,a)$ is unknown, our goal is to determine the optimal policy by utilizing the expert policies $\pi_{c},\pi_{e},\pi_{m}$. To accomplish this, we utilize the maximum casual entropy IRL method (ziebart2010modeling; ziebart2008maximum) to solve the following optimization problem:
@@
We study the policies obtained through RL, utilizing rewards learned through IRL on the most extensive set of reward functions, denoted as $\mathcal{R}$ in Eq.(8), which encompasses all functions mapping from $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ to $\mathbb{R}$. However, since using a large $\mathcal{R}$ can lead to overfitting in the IRL process, we employ a convex reward function regularizer (finn2016guided), denoted as $\psi$, to define the IRL procedure:
@@
(Theorem 2 of (syed2008apprenticeship)) If $\rho\in\Gamma$, then $\rho$ is the occupancy measure for $\pi_{\rho}(a|s)\triangleq\rho(s,a)/\sum_{a}^{\prime}\rho(s,a^{\prime})$, and $\pi_{\rho}$ is the only policy whose occupancy measure is $\rho$.
@@
Therefore, we can use minimax duality (millar1983minimax):
@@
ho2016generative present a cost regularizer, $\psi_{GA}$, that leads to an imitation learning algorithm, as outlined in Eq.(7), which aims to minimize the Jensen-Shannon divergence between the occupancy measures. Specifically, they convert a surrogate loss function, $\phi$, which is used for binary classification of state-action pairs drawn from the occupancy measures $\rho_{\pi}$ and $\rho_{\pi_{E}}$, into cost function regularizers $\phi$, such that $\phi^{*}(\rho_{\pi}-\rho_{\pi_{E}})$ represents the minimum expected risk, $R_{\phi}(\rho_{\pi},\rho_{\pi_{E}})$, for the function $\phi$ (ho2016generative):
@@
ho2016generative use the formula of surrogate loss function $\phi$: $\psi_{\phi}(c)=\sum_{\rho_{\pi_{E}}}g_{\phi}(c(s,a))$, where $g_{\phi}(x)=-x+\phi(-\phi^{-1}(-x))$, and $\phi$ is a strictly decreasing convex function (Appendix A.2 Proposition A.1 from (ho2016generative)). However, in our approach, we adopt reward function $r(s,a)$ instead of the cost function $c(s,a)$, thus we write in this form: $\psi_{\phi}(r)=\sum_{\rho_{\pi_{E}}}g_{\phi}(r(s,a))$, where $g_{\phi}(r(s,a))=x+\phi(-\phi^{-1}(x))$, and $\psi_{\phi}(r)$ is also closed, proper and convex (See proof in B.3).
@@
By using the logistic loss $\phi(\gamma)=\log(1+e^{-\gamma})$ (ho2016generative), we have $R_{\phi}(\rho_{\pi},\rho_{\pi_{E}})=\max\limits_{D\in(0,1)^{\mathcal{S}\times\mathcal{A}}}\sum_{s,a}\rho_{\pi}(s,a)\log D(s,a)+\rho_{\pi_{E}}(s,a)\log(1-D(s,a)),E=\{c,e,m\}$. Therefore, we obtain the final form of objective function as in Eq. (7).

### v2

@@ 5.1. Experiments Setup
Four real-world datasets are used: 1) PHEME (Zubiaga et al., 2016): This dataset collects posts from conversation threads during newsworthy events; 2) TWITTER (Ma et al., 2016): This dataset provides posts of events sourced from a fact-checking website, widely used for rumor detection studies; 3) Twitter-COVID-19 (Lin et al., 2022): This dataset gathers posts from conversation threads discussing events about the COVID-19 pandemic (Kar et al., 2020); 4) BEARD (Zeng and Gao, 2022): this dataset is for EARD task, focusing on covering early-stage posts relevant to various events. Each dataset offers a chronological collection of relevant posts for each event, annotated as either rumor or non-rumor.
@@
We compare with existing full-shot EARD methods under few-shot setup, including ERD (Zhou et al., 2019b) that uses a Deep Q-Network (DQN) to enforce the model to focus on early posts; CED (Song et al., 2019) that uses a fixed probability threshold to check if the prediction is credible for determining early detection point; HEARD (Zeng and Gao, 2022) that automatically determines a detection point at which predictions in future time steps remain unchanged using the neural Hawkes Process (Mei and Eisner, 2017). Given the strong capability of LLMs across NLP tasks, we also include Mistral (Jiang et al., 2023), Llama3 (Dubey et al., 2024) [^3] We refer to Mistral-Instruct-v0.3 as Mistral and Llama-3-Instruct as Llama3 for brevity., ChatGPT (OpenAI, 2024) as baselines. We incorporate these LLMs using two early detection strategies. Earliest Rumor Detection Strategy: The LLMs are prompted to make predictions using only the first post of each instance, following Miao et al. (2021). Preset Time Checkpoints Strategy: we use a set of preset time checkpoints at $\{1\text{h},6\text{h},12\text{h},24\text{h},36\text{h}\}$. All posts published before each checkpoint are provided to the LLMs for prediction. The timestamp of the first post serves as the starting point, and the time elapsed for each subsequent post is calculated relative to this initial timestamp. This strategy is commonly used in general rumor detection studies for evaluating early detection performance (Ma et al., 2016; Yu et al., 2017; Ma et al., 2017, 2018; Guo et al., 2018; Bian et al., 2020; Lin et al., 2023).
@@
We randomly sample 50 labeled training instances for trajectory generation and 100 unlabeled instances to serve as the environment for the agent, allowing the agent to dynamically explore and adapt its strategy in an online policy learning setting. We utilize the pretrained BERTweet model (Nguyen et al., 2020) to obtain embeddings of each state. Following Ho and Ermon (2016), we employ two hidden layers with a Tanh activation function for the policy, value, and discriminator networks. The coefficient of causal entropy $\lambda$ is set to 0.01. We set the expert ratios to $\alpha=0.7$, $\beta=0.15$, and $\gamma=0.15$ as discussed in §4.2.2.
@@
We report macro-F1 score for classification and use Early Rate (ER) (Song et al., 2019) to measure the proportion of posts used for detection. ER measures the proportion of posts used for detection: $ER=\frac{1}{|\mathbf{C}_{\text{test}}|}\sum_{C\in\mathbf{C}}\frac{i_{C}}{|C|}$, where $\mathbf{C}_{\text{test}}$ represents the test set, $i_{C}$ indicates the early detection decision made at the $i$-th post in instance $C$, and $|C|$ is the total number of posts in that instance. Lower ER means that the model can detect rumors earlier. For a fair comparison, we use the same prompt template for all models and run each experiment five times with different random seeds, reporting the mean and standard deviation for each metric. The prompt used for LLM prediction is “Analyze the given sequence of social media posts, determine if it is a rumor. Respond Yes or No only. Posts: [the list of observed posts in chronological order]”.
Table 3 presents hyper-parameters in our experiments. Following Ho and Ermon (2016), the policy, value, and discriminator networks are initialized randomly. We set the number of training steps to 200,000, with a batch size of 64 for the discriminator network and 4 for the policy and value networks.
@@ 5.2. Results and Analysis
First, as shown in Table 1, despite using only first post following the earliest rumor detection strategy (Miao et al., 2021), LLMs achieve competitive or superior macro-F1 scores across multiple datasets, compared to existing EARD methods. Specifically, ChatGPT outperforms all existing EARD methods by up to 33.3% and 10.3% on PHEME and Twitter-COVID-19, respectively. Even smaller LLMs like Mistral surpass all existing EARD methods on PHEME, and Llama3 outperforms ERD on TWITTER dataset. These results highlight the potential of LLMs used as detectors in EARD, particularly in data-scarce scenarios.
@@
Third, our method achieves overall better performance across both ER and F1, compared to existing EARD methods. For instance, in Table 1, on the TWITTER dataset, our framework enables Mistral to surpass all EARD baselines in F1 score while achieving a significantly lower ER (0.043) compared to ERD (1.000), CED (0.617), and HEARD (0.647). We also observe that ERD fails to make early decisions in few-shot setting, as shown by its high ER scores. This is likely due to its reliance on a manually designed reward function, where its policy trained by DQN is highly sensitive. Specifically, the reward function applies only a small penalty for continuation but imposes a large penalty for incorrect early termination (Zhou et al., 2019a), discouraging the model to stop early. This behavior aligns with previous findings (Zeng and Gao, 2022), and the problem is further exacerbated in the few-shot scenario, where data scarcity undermines the stability and effectiveness of reinforcement learning.
@@ 5.2.3. Analysis on How Generalizability Varies Across LLMs
As shown in Table 2 and discussed in § 5.2.1, our method generalizes better than the baseline EARD approaches to new event data. Trained on pre-COVID datasets (PHEME and TWITTER) and tested on Twitter-COVID-19, it consistently outperforms baselines in both F1 and ER. We further investigate how this generalizability varies across different base LLMs. As shown in Figure 4, our method dramatically boosts Llama3 by around 17% and 18% when trained on PHEME and TWITTER, respectively. This suggests that the agent can learn common expert decision patterns encoded in expert trajectories from other datasets, which benefits LLMs when applied to new event data. These gains are less pronounced with Mistral, which outperforms Llama3, and turn into a negative influence with ChatGPT, which initially shows the best performance on this COVID-related dataset. We conjecture that as the base model’s initial performance improves, the benefits of learned common behavior patterns become less useful compared to the information specific to the new event data. Thus, applying an agent trained on other datasets could lead to suboptimal decisions. One possible solution is to train the agent on a small number of new event data. As shown in Figure 4, even a few samples from the Twitter-COVID-19 dataset can improve performance across all base LLMs. Another promising direction is to model this problem as a Partially Observable Markov Decision Process (POMDP) (Sondik, 1971; Spaan, 2012), which is useful for capturing latent uncertainties and incomplete observations in sequential decision-making (Kaelbling et al., 1998) and may better capture the uncertainties in scenarios involving new event data. We leave this for future work.
@@ Appendix A Theoretical Analysis
The proof involves showing that the optimal policy in an imitation learning setting, typically obtained by first solving the Inverse Reinforcement Learning (IRL) problem to derive the optimal reward function $r^{*}$ and then applying a Reinforcement Learning (RL) algorithm, can be compressed into optimizing a $\psi$-regularized objective. Our method is based on generative adversarial imitation learning (Ho and Ermon, 2016), extending it to handle different types of expert trajectories in the context of LLM.
In the aforementioned equation, the objective is to find the saddle point of the minimax problem. Since the reward function $r(s,a)$ is unknown, our goal is to determine the optimal policy by utilizing the expert policies $\pi_{c},\pi_{e},\pi_{m}$. To accomplish this, we utilize the maximum casual entropy IRL method (Ziebart et al., 2010, 2008) to solve the following optimization problem:
@@
We study the policies obtained through RL, utilizing rewards learned through IRL on the most extensive set of reward functions, denoted as $\mathcal{R}$ in Eq.(8), which encompasses all functions mapping from $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ to $\mathbb{R}$. However, since using a large $\mathcal{R}$ can lead to overfitting in the IRL process, we employ a convex reward function regularizer (Finn et al., 2016), denoted as $\psi$, to define the IRL procedure:
@@
(Theorem 2 of (Syed et al., 2008)) If $\rho\in\Gamma$, then $\rho$ is the occupancy measure for $\pi_{\rho}(a|s)\triangleq\rho(s,a)/\sum_{a}^{\prime}\rho(s,a^{\prime})$, and $\pi_{\rho}$ is the only policy whose occupancy measure is $\rho$.
@@
Therefore, we can use minimax duality (Millar, 1983):
@@
Ho and Ermon (2016) present a cost regularizer, $\psi_{GA}$, that leads to an imitation learning algorithm, as outlined in Eq.(7), which aims to minimize the Jensen-Shannon divergence between the occupancy measures. Specifically, they convert a surrogate loss function, $\phi$, which is used for binary classification of state-action pairs drawn from the occupancy measures $\rho_{\pi}$ and $\rho_{\pi_{E}}$, into cost function regularizers $\phi$, such that $\phi^{*}(\rho_{\pi}-\rho_{\pi_{E}})$ represents the minimum expected risk, $R_{\phi}(\rho_{\pi},\rho_{\pi_{E}})$, for the function $\phi$ (Ho and Ermon, 2016):
@@
Ho and Ermon (2016) use the formula of surrogate loss function $\phi$: $\psi_{\phi}(c)=\sum_{\rho_{\pi_{E}}}g_{\phi}(c(s,a))$, where $g_{\phi}(x)=-x+\phi(-\phi^{-1}(-x))$, and $\phi$ is a strictly decreasing convex function (Appendix A.2 Proposition A.1 from (Ho and Ermon, 2016)). However, in our approach, we adopt reward function $r(s,a)$ instead of the cost function $c(s,a)$, thus we write in this form: $\psi_{\phi}(r)=\sum_{\rho_{\pi_{E}}}g_{\phi}(r(s,a))$, where $g_{\phi}(r(s,a))=x+\phi(-\phi^{-1}(x))$, and $\psi_{\phi}(r)$ is also closed, proper and convex (See proof in B.3).
@@
By using the logistic loss $\phi(\gamma)=\log(1+e^{-\gamma})$ (Ho and Ermon, 2016), we have $R_{\phi}(\rho_{\pi},\rho_{\pi_{E}})=\max\limits_{D\in(0,1)^{\mathcal{S}\times\mathcal{A}}}\sum_{s,a}\rho_{\pi}(s,a)\log D(s,a)+\rho_{\pi_{E}}(s,a)\log(1-D(s,a)),E=\{c,e,m\}$. Therefore, we obtain the final form of objective function as in Eq. (7).